The amount of money in an account earning compound interest is modelled by the equation:
P(n) = Arn – 1 where P(n) is the amount in the account at year n, A is the amount invested, and r = 1 + i,
and i expressed as a percent is the effective rate of interest. If a man invests $1500 in year 1, and earns
4% interest, how much money will he have in year 30?
a. $4677.98 c. $6482.91
b. $5059.70 d. $4865.10
Answer #1
P(n) =A*(1+i)^n
Here,
A =$1,500
I =4% and 1+I =1+0.4 =1.04
n =30
$1,500*(1.04)^30
$1,500*3.2434 =$4,865.10
So Option D is the answer